OpenStudy (anonymous):

I need to find all vectors v that are orthogonal to u. Where u = (2,-1,1)

jimthompson5910 (jim_thompson5910):

Let v = (x, y , z) where v is a vector and x,y,z are scalars

jimthompson5910 (jim_thompson5910):

if vector v is orthogonal to vector u, then the dot product of the two vectors should be 0 v dot u = 0 (x,y,z) dot (2,-1,1) = 0 2*x + (-1)*y + 1*z = 0 2x - y + z = 0

jimthompson5910 (jim_thompson5910):

you can then solve that for z to get 2x - y + z = 0 -y + z = -2x z = -2x + y

OpenStudy (anonymous):

okay, thank you that makes sense, but what do i do from here? do I also need to solve for x and y?

jimthompson5910 (jim_thompson5910):

no, you now let x = s y = t where x,y,s,t are all scalars. This allows you to parameterize things so v = (x,y,z) becomes v = (s,t,-2s+t) that's the general vector v that is orthogonal to vector u. To generate a specific vector v (that has numbers instead of variables, you can pick any ordered pair (s,t) and plug them in)

OpenStudy (anonymous):

Oh! thank you! that makes sense. That helped a lot thank you for your help!

jimthompson5910 (jim_thompson5910):

you're welcome

jimthompson5910 (jim_thompson5910):

I'm glad it's making sense now